First calculate the Income of your family by using the following formula:

Family Income = Last Two digit of your enrolment number * Parents + Number of siblings

* 500


1
Expert's answer
2021-04-20T07:24:47-0400

Family Income = Last Two digit of your enrollment number "\times" Parents + Number of siblings

"\times" 500 = "53\times 2+2\times500=\$1106"